For the following problems solve the IVP using Laplace Method - be careful of initial conditions and coefficients which change in each problem:

a. y" + 5y' + 6y = 5e-5t ; y'(0) = 0, y(0) = 0

b. y' + 6y = t ; y(0) = 0
Answer #1

Page 3 y Co) = 0 (6 446y=t Apply Laplace transform on both side yt by=t = Llyt64) = [ct = LY)+6 Lly) = LCt) By wring Ley)